OpenStudy (anonymous):

Find the Domain and Range of the Function f(x,y) =sqrt(x^2 + y^2 + 4).

OpenStudy (jmark):

Given Function f(x,y) =sqrt(x^2 + y^2 + 4) the domain of the function =R^2 {R is real valued } and rang of the function ={f belongs to R:f>=2} See also Algebra Word Problems http://goo.gl/sA1qjA